OpenStudy (anonymous):

Determine if the series abs converges, conditionally converges, or diverges. n^n /n!

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

I don't know what test to use for convergence. For abs. conv. I used ratio test and well I got the limit as n approaches infinity of (n+-)^n / n^n

OpenStudy (anonymous):

I'm not sure if its infinity or 1. There's a shortcut for when the limit approaches infinity, so that when the power on the top is the same as the bottom you take the coefficients. But is that just for polynomial functions?

OpenStudy (anonymous):

Now you have \[\frac{(n+1)^{n}}{n^{n}}=\left( \frac{n+1}{n} \right)^{n}=\left( 1+\frac{1}{n} \right)^{n}\] Oh that looks familar

OpenStudy (anonymous):

e

OpenStudy (anonymous):

Yup. So what conclusion do you reach about the series

OpenStudy (anonymous):

Dang, I forgot about that :/ thanks! And it diverges

OpenStudy (anonymous):

Yup. Sure thing

OpenStudy (anonymous):

Do you have a suggestion on what test to use to see if it conditionally converges?

OpenStudy (anonymous):

Its usually alternating series that do that, right. There is a test for that

OpenStudy (anonymous):

I didn't think so since its not alternating sign. It's positive

OpenStudy (anonymous):

Ah sorry. to see if a series converges conditionally see if the absolute value also converges. If it doesn't then you have conditional convergence

OpenStudy (anonymous):

But don't you have to see if it converges without taking the absolute. Because they both can diverge

OpenStudy (anonymous):

Yes. I mean after you confirm that the original series converges, check if the absolute does as well to check if its conditional

OpenStudy (anonymous):

That's my problem, I don't know what test to use to see if original series converges

OpenStudy (anonymous):

What is the series?

OpenStudy (anonymous):

n^n /n!